# Central galaxy alignments

### v1

@@ 5 Environmental effects
Figure 6: Dependence on mass the angles ($\theta_{int}$) (left), ($\theta_{sat}$) (centre), and ($\theta_{ch}$) (right) for central galaxies located in filaments of varying densities. Darker lines represent the highest density quartile, while lighter lines correspond to the lowest. In all panels, solid lines represent the major axis, while dotted lines represent the minor axis. The lower panels show the ratio between the results at different densities.
@@
Having established the relationship between the studied angles and halo mass, we now will analyze how the large-scale environment and its characteristics influence these alignments in filaments, clusters, cluster outskirts, and others. Given that clusters occupy uniquely the high mass-end of the mass distribution, the trends for galaxies in these structures will be dominated by the mass effect described above. Regardless, we will include this information for the sake of completeness. In contrast, the remaining environments are populated by objects within a similar and overlapping mass range, as illustrated in Figure 1, thus enabling the exploration of any specific trends arising from environmental effects while keeping the mass fixed. To begin with, we analyze the alignments taking into account the axes defined using dark matter particles. This builds on the results derived from the black curves presented in the previous section.
In the left panel of Figure 4, we illustrate how $\theta_{int}$ varies as a function of mass across different environments. At a glance, we can see that in all environments the same trend with mass is observed as in the overall sample of the previous section/figure. Along the major axis (solid lines), we observe that around $10^{13}\,M_{\odot}$ stellar material and dark matter exhibit lower alignment in filaments than others, as the $\theta_{int}$ values for galaxies in filaments are  $\sim\,4^{\circ}$ higher than in others. Note that the cluster outskirts shows noisy results, mainly because there are only 2122 halos in this environment. In contrast, galaxies in others demonstrate the lowest degree of alignment at lower masses. Additionally, within the mass range from approximately $10^{12}$ to $10^{13}\,M_{\odot}$, central galaxies in filaments are less aligned compared to those in other environments. Although the above results are statistically significant, the differences found are minimal across the different environments compared to the trends with mass, which are the dominant ones driving the alignments.
@@ 5.1 Variations with Filament density
In Figure 6, we analyze the behaviour of the angles $\theta_{int}$, $\theta_{sat}$, and $\theta_{ch}$ by dividing the filaments based on their density. Two samples were taken using the first and fourth quartiles of the filament densities, corresponding to 24.5 and 82.7 $Mpc^{-3}$, respectively. Note that, to calculate this density, we considered all galaxies with stellar mass greater than $10^{8.5}M_{\odot}$. Darker lines represent higher-density filaments, while lighter lines indicate lower-density ones. In the left panel, we examine the alignment between dark matter and stars ($\theta_{int}$). For denser filaments, the stellar component is less aligned with its dark matter halo along the major axis, particularly at higher masses. In contrast, the minor axes exhibit the opposite trend, with greater alignment in denser environments. This is clearly shown in the lower panel, which quantifies the ratio of angles between low- and high-density filaments. For the major axis (solid line), we observe differences of up to 20% at high masses, while the minor axis (dashed line) shows smaller differences of around 10% at low masses. The middle and right panels present $\theta_{sat}$ and ($\theta_{ch}$), respectively. Both panels show that for both minor and major axes, there is a greater alignment in less dense filaments. In summary, the only exception to this trend is the alignment $\theta_{int}$ for the minor axis at lower masses, which behaves differently than the other alignments. Regarding the ratio of angles between low- and high-density filaments for $\theta_{sat}$, our analysis demonstrates that only the major axis is responsive to filament density, showing consistent differences of approximately 10% throughout the mass range. Furthermore, the ratio for $\theta_{ch}$ remains similar for both axes and increases with mass, spanning from very low values to greater than 20%. These results underscore the strong effect of different filament environments on the alignment of central galaxies with respect to the three different references explored in this paper.
The results indicate that in the outermost regions of the halo traced by $\theta_{sat}$ and $\theta_{ch}$, the denser filaments generate a misalignment with respect to the mayor and minor axes of the central galaxy. On the other hand, in the innermost regions (traced by $\theta_{int}$), dark matter and stars tend to align with the minor axis of the galaxy.
@@ 5.2 Variations with local density
At this stage, a key question emerges: are the filaments (as physical structures) directly influencing the alignments, or are we simply measuring local density effects? In other words, are the filaments or the local environment responsible for affecting the alignments? This question arose while analyzing the differences in observed alignments at the extremes of the filament density distribution. Therefore, it is important to examine how much these samples differ in local density.
To address this issue, we estimate local density by calculate the density of dark matter within a sphere of radius 5 Mpc centred around each BGG. We compare the local density distributions of central galaxies found in low- and high-density filaments, as illustrated in Figure 7. We present the results across four mass ranges to avoid potential biases associated with mass variability. Due to the significant differences in local density observed between the two samples, we found that adjusting for local density resulted in alignment differences that closely resembled those shown in Figure 6, which classifies galaxies according to filament density (we do not display this figure here because it does not provide relevant information).
We take a step further and focus now on the differences between centrals in filaments and those in others, i.e. the most under-dense cosmic environments. Figure 8 illustrates the density distribution for those galaxies classified as belonging to filaments and others. As can be expected, the distributions differ significantly, with the filaments being typically much denser. In order to determine whether the presence of filaments has any effect on the alignments beyond local density effects, we make the two samples comparable by selecting a subsample of BGGs from the others that match the density distribution of the filaments sample. This is illustrated by the dotted lines in Figure 8.
@@ 6 SDSS
Figure 10: SDSS measurements of the alignments of the major axis of the central galaxies with the location of the satellite galaxies ($\varphi_{sat}$) and of the major axis of the central galaxies with the major axis of the group to which they belong, measured with the member galaxies ($\varphi_{ch}$). The upper panels show the total sample and its dependence on mass, while the lower panels show those corresponding to different environments of the cosmic web.
@@
We first note that, despite the projection, we recover the trends detected in the simulations, as the alignment of the satellites $\varphi_{sat}$ (shown in the upper left panel) shows a strong dependence on the mass. Similar to the simulations, no alignment is observed in low-mass systems; however, alignment significantly increases as mass increases (for reference, see measurements indicated by the yellow lines in Figure 3). It is remarkable to find this dependence in the observations and that it agrees with the simulations. Indeed this result supports the reliability of the abundance matching procedure, which assigns masses in the group catalogue based solely on luminosity ranking and aims to provide a proxy for the mass, rather than an exact measurement.
In the upper right panel of figure 10, we show the projected alignment of the BGG shape with respect to the halo shape, $\varphi_{ch}$. Once again, the bahaviour of $\varphi_{ch}$ as a function of mass resembles that of the simulations: for low masses, the alignment is undetectable, whereas for high masses, it increases. It should be emphasised that, besides the uncertainties in determining the mass, the shapes are estimated with very few members and, therefore, the direction of the major axis of the group has significant uncertainty, unlike the simulations where the shape is determined with greater precision using numerous dark matter particles.
@@ 7 Summary and conclusions
We analyse how three different angles characterising the alignment of central galaxies with their multi-scale environment to understand their relation to the structure on different scales. First, we study the alignment between the stellar and dark matter components of the central galaxy ($\theta_{int}$), which provides information on the internal alignment. Second, we examine the angle between the satellite galaxies and the shape axes of the central galaxy ($\theta_{sat}$), allowing us to characterise the distribution of stellar material within the halo in which the group resides. Finally, we investigate the alignment between the shape axis of the central galaxies and that of the group’s host halo ( $\theta_{ch}$), providing information on the connection between galaxy and structure on larger scales These angles are illustrated in 2, which provides a schematic representation of their definitions. These analyses let us explore how the surroundings and multi-scale physical processes influence the angles characterising the central galaxies.
@@
First, we performed our analysis using the TNG300 hydrodynamical simulation. We explored how each angle depended on the halo mass. We found that both $\theta_{\text{int}}$ and $\theta_{\text{ch}}$—measured using the shape axes of the central galaxy defined by either the dark matter or stellar distributions— exhibit a clear trend of increasing with mass (see Figure 3). The first one suggests that, internally, the stellar component becomes more tightly coupled with dark matter as the halo mass increases. Similarly, from $\theta_{\text{ch}}$ we can infer that the dark matter and stellar components of the central galaxy show a stronger alignment with the dark matter of the host halo as mass grows. In contrast, $\theta_{\text{sat}}$ does not demonstrate a clear dependence on mass when measured using the dark matter axes. When we consider the axes defined by the stellar distribution, a clear increasing trend with mass emerges. The distribution of stars within central galaxies depends on mass, and this dependence manifests itself as varying alignment with increasing mass. Additionally, in all cases where we measure the angles using the axes defined by stellar distribution, we find that it is systematically weaker than when measured using the dark matter axes. This suggests that while the stellar structure somewhat follows the dark matter, its degree of alignment with its surroundings is less pronounced. This could mean that the astrophysical processes affecting stars are more complex than those influencing dark matter, which is reflected in their alignments.
@@
Another analysis we conducted was the study of the variation of alignments as a function of mass across different environments of the cosmic web (see Figure 4). We classified these environments into four categories: clusters, filaments, cluster outskirts, and a category labelled others, which includes voids and walls. As expected, the most massive central galaxies —also exhibiting the strongest alignments— are by definition in clusters, while the other three environments host central galaxies of lower masses. We found significant differences between environments, suggesting that alignments depend on large-scale structure. Since we compare galaxies at fixed mass, the detected variations reflect the impact of the environment beyond the effect of the host halo mass. While the mass of the host halo largely influences alignments, the location in the most hostile environments of the cosmic web appears to disrupt these alignments at a fixed mass. Furthermore, since our previous analyses were based on dark matter, we investigated whether these differences persisted when measuring the principal axes of central galaxies using their stellar component (Figure 5). We observed that the differences decreased compared to those found with dark matter, indicating that environmental influence is more pronounced in the dark matter structure than in the stellar distribution.
@@
To contrast the results of the TNG300 simulations, we performed a similar analysis using SDSS data. First, we examine the alignments as a function of mass (see Figure 10) and find a clear dependence, similar to what we observed in the simulations. This finding is particularly significant given the numerous observational biases inherent in SDSS data that could potentially obscure the signal, unlike in simulations where such biases are absent. The consistency between the two analyses underscores the reliability of group identification and mass assignments in the observational data. Additionally, it suggests that the physical effects that influence galaxy alignments in the real universe are reflected in the simulations. Secondly, we investigated whether the differences observed between environments in the simulations were also present in the observational data, focusing on three environments: clusters, filaments and others. However, we did not detect significant differences, likely due to observational biases that blur the effects identified in simulations. Nevertheless, it would be interesting to explore this further using alternative filament and void identification methods applied to future observational data.
@@
In conclusion, this study has deepened our understanding of the alignments of central galaxies concerning their environment, both internally and on larger scales, using hydrodynamical simulations and observational data. The results confirm that the halo mass, and environment (both local and, to a lesser extent, large-scale), play a crucial role in determining these alignments. However, open questions remain about how other factors, such as the processes of galaxy formation and evolution, might influence these patterns. Future research could explore in more detail the impact of the accretion history of galaxies or the relationship between galaxy morphology and their alignment in different environments. Additionally, it would be beneficial to validate these findings through larger volume simulations, which would enhance statistical analysis. Exploring new methods for identifying structures in observational data would also help refine our understanding of the connection between galaxies and the cosmic web. This work is particularly timely, as new data from upcoming spectroscopic surveys such as DESI (Dark Energy Spectroscopic Instrument), Euclid, and PFS (Prime Focus Spectrograph) will soon be available. These surveys will provide unprecedented statistics and resolution, enabling rigorous tests of our predictions and offering new insights into galaxy alignments within the cosmic web.
@@ Acknowledgements.
FR, MM and AVMC thanks the support by Agencia Nacional de Promoción Científica y Tecnológica, the Consejo Nacional de Investigaciones Científicas y Técnicas (CONICET, Argentina) and the Secretaría de Ciencia y Tecnología de la Universidad Nacional de Córdoba (SeCyT-UNC, Argentina). FR would like to acknowledge support from the ICTP through the Junior Associates Programme 2023-2028. DGE acknowledges Volker Springel for the financial support that facilitated the research visit during which part of this work was conducted. MCA acknowledges support from ANID BASAL project FB210003. ADMD thanks Fondecyt for financial support through the Fondecyt Regular 2021 grant 1210612.

### v2

@@ 5 Environmental effects
Figure 6: Dependence on halo mass the angles $\theta_{int}$ (left), $\theta_{sat}$ (centre), and $\theta_{ch}$ (right) for central galaxies located in filaments of varying densities. Darker lines represent the highest density quartile, while lighter lines correspond to the lowest. In all panels, solid lines represent the major axis, while dotted lines represent the minor axis. The lower panels show the ratio between the results at different densities.
@@
Having established the relationship between the studied angles and halo mass, we now will analyze how the large-scale environment and its characteristics influence these alignments in filaments, clusters, cluster outskirts, and others. Given that clusters occupy uniquely the high mass-end of the halo mass distribution, the trends for galaxies in these structures will be dominated by the mass effect described above. Regardless, we will include this information for the sake of completeness. In contrast, the remaining environments are populated by objects within a similar and overlapping halo mass range, as illustrated in Figure 1, thus enabling the exploration of any specific trends arising from environmental effects while keeping the halo mass fixed. To begin with, we analyze the alignments taking into account the axes defined using dark matter particles. This builds on the results derived from the black curves presented in the Figure 3.
In the left panel of Figure 4, we illustrate how $\theta_{int}$ varies as a function of halo mass across different environments. At a glance, we can see that in all environments the same trend with halo mass is observed as in the overall sample of the previous section/figure. Along the major axis (solid lines), we observe that around $10^{13}\,M_{\odot}$ stellar material and dark matter exhibit lower alignment in filaments than others, as the $\theta_{int}$ values for galaxies in filaments are  $\sim\,4^{\circ}$ higher than in others. Note that the cluster outskirts shows noisy results, mainly because there are only 2122 halos in this environment. In contrast, galaxies in others demonstrate the lowest degree of alignment at lower masses. Additionally, within the halo mass range from approximately $10^{12}$ to $10^{13}\,M_{\odot}$, central galaxies in filaments are less aligned compared to those in other environments. Although the above results are statistically significant, the differences found are minimal across the different environments compared to the trends with mass, which are the dominant ones driving the alignments.
@@ 5.1 Variations with Filament density
In Figure 6, we analyze the behaviour of the angles $\theta_{int}$, $\theta_{sat}$, and $\theta_{ch}$ by dividing the filaments based on their density. Two samples were taken using the first and fourth quartiles of the filament densities, corresponding to 24.5 and 82.7 $Mpc^{-3}$, respectively. Note that, to calculate this density, we considered all galaxies with stellar mass greater than $10^{8.5}M_{\odot}$. Darker lines represent higher-density filaments, while lighter lines indicate lower-density ones. In the left panel, we examine the alignment between dark matter and stars ($\theta_{int}$). For denser filaments, the stellar component is less aligned with its dark matter halo along the major axis, particularly at higher masses. In contrast, the minor axes exhibit the opposite trend, with greater alignment in denser environments. This is clearly shown in the lower panel, which quantifies the ratio of angles between low- and high-density filaments, $\theta^{low}/\theta^{high}$. For the major axis (solid line), we observe differences of up to 20% at high masses, while the minor axis (dashed line) shows smaller differences of around 10% at low masses. The middle and right panels present $\theta_{sat}$ and $\theta_{ch}$, respectively. Both panels show that for both minor and major axes, there is a greater alignment in less dense filaments. In summary, the only exception to this trend is the alignment $\theta_{int}$ for the minor axis at lower masses, which behaves differently than the other alignments. Regarding the ratio of angles between low- and high-density filaments for $\theta_{sat}$, our analysis demonstrates that only the major axis is responsive to filament density, showing consistent differences of approximately 10% throughout the halo mass range. Furthermore, the ratio for $\theta_{ch}$ remains similar for both axes and increases with mass, spanning from very low values to greater than 20%. These results underscore the strong effect of different filament environments on the alignment of central galaxies with respect to the three different reference scales explored in this paper.
The results indicate that in the outermost regions of the halo traced by $\theta_{sat}$ and $\theta_{ch}$, the denser filaments generate a misalignment with respect to the major and minor axes of the central galaxy. On the other hand, in the innermost regions (traced by $\theta_{int}$), dark matter and stars tend to align with the minor axis of the galaxy.
@@ 5.2 Variations with local density
At this stage, a key question emerges: are the filaments (as physical structures) directly influencing the alignments, or are we simply measuring local density effects? In other words, are the filaments or the local environment responsible for affecting the alignments? Therefore, it is important to examine how much these samples differ in local density.
To address this issue, we estimate local density by calculating the density of dark matter within a sphere of radius 5 Mpc centred around each BGG. We compare the local density distributions of central galaxies found in low- and high-density filaments, as illustrated in Figure 7. We present the results across four halo mass ranges to avoid potential biases associated with the mass variability. Due to the significant differences in local density observed between the two samples, we found that adjusting for local density resulted in alignment differences that closely resembled those shown in Figure 6, which classifies galaxies according to filament density (we do not display this figure here because it does not provide relevant information).
We take a step further and focus now on the differences between centrals in filaments and those in others, i.e. the most under-dense cosmic environments. Figure 8 illustrates the local density distribution for those galaxies classified as belonging to filaments and others. As can be expected, the distributions differ significantly, with the filaments being typically much denser. In order to determine whether the presence of filaments has any effect on the alignments beyond local density effects, we make the two samples comparable by selecting a subsample of BGGs from the others that match the density distribution of the filaments sample. This is illustrated by the dotted lines in Figure 8.
@@ 6 SDSS
Figure 10: SDSS measurements of the alignments of the major axis of the central galaxies with the location of the satellite galaxies ($\varphi_{sat}$) and of the major axis of the central galaxies with the major axis of the group to which they belong, measured with the member galaxies ($\varphi_{ch}$). The upper panels show the total sample and its dependence on mass, while the lower panels show those corresponding to different environments of the cosmic web. In all panels, error bars represent the standard deviation of the mean.
@@
We first note that, despite the projection, we recover the trends detected in the simulations, as the alignment of the satellites $\varphi_{sat}$ (shown in the upper left panel) shows a strong dependence on the mass. Similar to the simulations, no alignment is observed in low-mass systems; however, alignment significantly increases as halo mass increases (for reference, see measurements indicated by the yellow lines in Figure 3). It is remarkable to find this dependence in the observations and that it agrees with the simulations. Indeed this result supports the reliability of the abundance matching procedure, which assigns masses in the group catalogue based solely on luminosity ranking and aims to provide a proxy for the mass, rather than an exact measurement.
In the upper right panel of figure 10, we show the projected alignment of the BGG shape with respect to the halo shape, $\varphi_{ch}$. Once again, the behaviour of $\varphi_{ch}$ as a function of halo mass resembles that of the simulations: for low masses, the alignment is undetectable, whereas for high masses, it increases. It should be emphasised that, besides the uncertainties in determining the mass, the shapes are estimated with very few members and, therefore, the direction of the major axis of the group has significant uncertainty, unlike the simulations where the shape is determined with greater precision using numerous dark matter particles.
@@ 7 Summary and conclusions
We analyse how three different angles characterise the alignment of central galaxies with their multi-scale environment to understand their relation to the structure on different scales. First, we study the alignment between the stellar and dark matter components of the central galaxy ($\theta_{int}$), which provides information on the internal alignment. Second, we examine the angle between the satellite galaxies and the shape axes of the central galaxy ($\theta_{sat}$), allowing us to characterise the distribution of stellar material within the halo in which the group resides. Finally, we investigate the alignment between the shape axis of the central galaxies and that of the group’s host halo ($\theta_{ch}$), providing information on the connection between galaxy and structure on larger scales These angles are illustrated in Figure 2, which provides a schematic representation of their definitions. These analyses let us explore how the surroundings and multi-scale physical processes influence the angles characterising the central galaxies.
@@
First, we performed our analysis using the TNG300 hydrodynamical simulation. We explored how each angle depended on the halo mass. We found that both $\theta_{\text{int}}$ and $\theta_{\text{ch}}$—measured using the shape axes of the central galaxy defined by either the dark matter or stellar distributions— exhibit a clear trend of decreasing with halo mass (see Figure 3). The first one suggests that, internally, the stellar component becomes more tightly coupled with dark matter as the halo mass increases. Similarly, from $\theta_{\text{ch}}$ we can infer that the dark matter and stellar components of the central galaxy show a stronger alignment with the dark matter of the host halo as mass grows. In contrast, $\theta_{\text{sat}}$ does not demonstrate a clear dependence on mass when measured using the dark matter axes. When we consider the axes defined by the stellar distribution, a clear increasing trend with halo mass emerges. However, in all cases where we measure the angles using the axes defined by stellar distribution, we find that it is systematically weaker than when measured using the dark matter axes. This suggests that while the stellar structure somewhat follows the dark matter, its degree of alignment with its surroundings is less pronounced. This could mean that the astrophysical processes affecting stars are more complex than those influencing dark matter, which is reflected in their alignments. These results expand on previous works, such as Valenzuela et al. (2024), where the authors study central galaxy shapes in the Magneticum simulation and conclude that stellar and dark matter shapes are correlated at internal scales, but the inner halo appears to be decoupled from the outer halo. We find that the strength of both the internal stellar–subhalo and the subhalo–host halo alignment are strongly coupled in the more massive halos, and the large-scale structure environment impacts more on the subhalo–host halo alignment than on the internal alignment.
@@
Another analysis we conducted was the study of the variation of alignments as a function of halo mass across different environments of the cosmic web (see Figure 4). We classified these environments into four categories: clusters, filaments, cluster outskirts, and a category labelled others, which includes voids and walls. As expected, the most massive central galaxies —also exhibiting the strongest alignments— are by definition in clusters, while the other three environments host central galaxies of lower masses. We found significant differences between environments, suggesting that alignments depend on large-scale structure. Since we compare galaxies at fixed mass, the detected variations reflect the impact of the environment beyond the effect of the host halo mass. While the mass of the host halo largely influences alignments, the location in the most hostile environments of the cosmic web appears to disrupt these alignments at a fixed halo mass. Furthermore, since our previous analyses were based on dark matter, we investigated whether these differences persisted when measuring the principal axes of central galaxies using their stellar component (Figure 5). We observed that the differences decreased compared to those found with dark matter, indicating that environmental influence is more pronounced in the dark matter structure than in the stellar distribution.
@@
To contrast the results of the TNG300 simulations, we performed a similar analysis using SDSS data. First, we examine the alignments as a function of halo mass (see Figure 10) and find a clear dependence, similar to what we observed in the simulations. This finding is particularly significant given the numerous observational biases inherent in SDSS data that could potentially obscure the signal, unlike in simulations where such biases are absent. The consistency between the two analyses underscores the reliability of group identification and halo mass assignments in the observational data. Additionally, it suggests that the physical effects that influence galaxy alignments in the real universe are reflected in the simulations. Secondly, we investigated whether the differences observed between environments in the simulations were also present in the observational data, focusing on three environments: clusters, filaments and others. However, we did not detect significant differences, likely due to observational biases that blur the effects identified in simulations. Nevertheless, it would be interesting to explore this further using alternative filament and void identification methods applied to future observational data.
@@
In conclusion, this study has deepened our understanding of the alignments of central galaxies concerning their environment, both internally and on larger scales, using hydrodynamical simulations and observational data. The results confirm that the halo mass, and environment (both local and, to a lesser extent, large-scale), play a crucial role in determining these alignments. However, open questions remain about how other factors, such as the processes of galaxy formation and evolution, might influence these patterns. For example, Xu et al. (2023b) analyzed the alignment between the principal axes of the stellar component and the host halo as a function of mass, and showed that fixing the ex situ stellar mass fraction almost eliminates the alignment’s dependence on halo mass, meaning that mass accretion processes strongly influence the alignments. Also, other studies using the SAMI and MaNGA surveys have established that there exists a relationship between the stellar mass and the alignment of galaxies’ angular momentum with filaments, and that some evolutionary processes such as bulge growth, AGN activity and mergers play an important role in producing the so-called spin flip effect (Welker et al., 2020; Kraljic et al., 2021; Barsanti et al., 2022, 2023, 2025). Given that galaxy shape orientation can be related to their stellar angular momentum, there may be a connection between the spin–filament alignment and the axis alignments studied in this work. Nonetheless, caution is needed when making comparisons, as these are different types of alignments. Therefore, future research projects could explore in more detail the impact of the accretion history of galaxies or the relationship between galaxy morphology and their alignment in different environments. Additionally, it would be beneficial to validate these findings through larger volume simulations like Flamingo (Schaye et al., 2023) or MillenniumTNG (Pakmor et al., 2023), which would enhance statistical analysis. Exploring new methods for identifying structures in observational data would also help refine our understanding of the connection between galaxies and the cosmic web. This work is particularly timely, as new data from upcoming spectroscopic surveys such as DESI (Dark Energy Spectroscopic Instrument), Euclid, and PFS (Prime Focus Spectrograph) will soon be available. These surveys will provide unprecedented statistics and resolution, enabling rigorous tests of our predictions and offering new insights into galaxy alignments within the cosmic web.
@@ Acknowledgements.
The authors wish to thank the anonymous referee for her/his report that helps us to improve this manuscript. FR, MM and AVMC thanks the support by Agencia Nacional de Promoción Científica y Tecnológica, the Consejo Nacional de Investigaciones Científicas y Técnicas (CONICET, Argentina) and the Secretaría de Ciencia y Tecnología de la Universidad Nacional de Córdoba (SeCyT-UNC, Argentina). FR and ADMD thank the ICTP for their hospitality and financial support through the Junior Associates Programme 2023–2028 and Senior Associates Programme 2022–2027, respectively. DGE acknowledges Volker Springel for the financial support that facilitated the research visit during which part of this work was conducted. MCA acknowledges support from ANID BASAL project FB210003. ADMD thanks Fondecyt for financial support through the Fondecyt Regular 2021 grant 1210612.